# Decentralized and Uncoordinated Learning of Stable Matchings: A Game-Theoretic Approach

### v1

@@ III-B Performance of Algorithm 1 for Hierarchical Markets
$Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w}$ $=\sum_{\tau=1}^{t-1}\big{(}v_{m_{k}w_{k}}(\hat{X}^{\tau})-v_{m_{k}w}(\hat{X}^{\tau})\big{)}$
$\geq\sum_{\tau=1}^{t-1}\Big{(}\mu_{m_{k}w_{k}}\prod_{\ell=1}^{k-1}(1-\hat{X}^{\tau}_{m_{\ell}w_{k}})-\mu_{m_{k}w}\Big{)}$ (37)
$\geq\sum_{\tau=1}^{t-1}\Big{(}\mu_{m_{k}w_{k}}\big{(}1-\sum_{\ell=1}^{k-1}\hat{X}^{\tau}_{m_{\ell}w_{k}}\big{)}-\mu_{m_{k}w}\Big{)}$ (38)
$\geq\sum_{\tau=1}^{t-1}\Big{(}(\mu_{m_{k}w_{k}}-\mu_{m_{k}w})-\sum_{\ell=1}^{k-1}\big{(}\frac{\gamma^{t}}{n}+\frac{(1-\gamma^{t})e^{-\eta^{t}a_{\ell}t}}{1+(n-1)e^{-\eta^{t}a_{\ell}t}}\big{)}\Big{)}$ (39)
$\geq\sum_{\tau=1}^{t-1}\Big{(}\Delta-\gamma^{t}-\sum_{\ell=1}^{k-1}e^{-\eta^{t}a_{\ell}t}\Big{)}$ (40)
$\geq\sum_{\tau=1}^{t-1}\Big{(}\Delta-\gamma^{t}-(n-1)e^{-\eta^{t}a_{k-1}t}\Big{)}.$ (41)
Thus, for any integer $t\geq t_{k-1}$ that satisfies $e^{-\eta^{t}a_{k-1}t}\leq\gamma^{t}$ and $\frac{t-2}{\sum_{\tau=1}^{t-1}\gamma^{\tau}}\geq\frac{2n}{\Delta}$, we have
$Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w}\geq\sum_{\tau=1}^{t-1}\Big{(}\Delta-n\gamma^{t}\Big{)}=\Delta(t-1)-n\sum_{\tau=1}^{t-1}\gamma^{\tau}\geq\frac{\Delta}{2}t.$ (42)
@@
$t_{k}:=\min\Big{\{}t\geq t_{k-1}|\ \ \frac{t\eta^{t}}{\ln(1/\gamma^{t})}\geq\frac{1}{a_{k-1}},\ \frac{t-2}{\sum_{\tau=1}^{t-1}\gamma^{\tau}}\geq\max\{\frac{4n}{\mu_{m_{k}w_{k}}},\frac{2n}{\Delta}\}\Big{\}},$ (43)
@@
$\frac{X^{t}_{m_{k}w_{k}}}{X^{t}_{m_{k}w}}=e^{\eta^{t}(\hat{L}^{t-1}_{m_{k}w_{k}}-\hat{L}^{t-1}_{m_{k}w})}\geq e^{\eta^{t}(Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w})-2c\sqrt{t}}\geq e^{\eta^{t}2a_{k}t-2c\sqrt{t}}\geq e^{\eta^{t}2a_{k}t-\eta^{t}a_{k}t}=e^{\eta^{t}a_{k}t},$
@@
Finally, using the stepsize and mixing parameters $\eta^{t}=\frac{1}{\sqrt{t}}$ and $\gamma^{t}=M\frac{\log t}{t}$ in relation (43), one can obtain an explicit upper bound for $t_{n}$ as
$t_{n}\leq\min\Big{\{}t\in\mathbb{Z}_{+}|\ \ \frac{\sqrt{t}}{\log(t/M\log t)}\geq\frac{1}{a_{n-1}},\ \frac{t-2}{M\log^{2}t}\geq\max\{\frac{4n}{\mu_{m_{n}w_{n}}},\frac{2n}{\Delta}\}\Big{\}},$
for which it suffices to take $t_{n}=\tilde{O}\big{(}\frac{n^{2}}{c^{2}}\log\frac{1}{\delta}\big{)}$, where $\tilde{O}(\cdot)$ hides the logarithmic factors.  Q.E.D.
@@
Assume that each man follows Algorithm 1 with $\eta^{t}=\frac{1}{\sqrt{t}}$ and $\gamma^{t}=M\frac{\log t}{t}$, where $M=\frac{4n}{c}\log T$ and $c=\frac{1}{8}\min_{k\in[n]}\{\Delta,\mu_{m_{k}w_{k}}\}$. Then, the expected regret of Algorithm 1 in hierarchical matching markets is at most $R(T)=O\Big{(}\frac{n^{2}}{c}\log^{3}T\Big{)}$.
@@
Conditioned on the event $\Omega$, which according to Lemma 2 it occurs with probability at least $1-\delta$, the expected regret can be upper-bounded as
$\mathbb{E}[R(T)|\Omega]$ $=\sum_{t=1}^{T}\mathbb{E}[\boldsymbol{1}_{\{\exists k:\alpha^{t}_{m_{k}}\neq w_{k}\}}|\Omega]\leq\sum_{t=1}^{T}\sum_{k=1}^{n}\mathbb{E}[\boldsymbol{1}_{\{\alpha^{t}_{m_{k}}\neq w_{k}\}}|\Omega]$
$\leq\sum_{t=1}^{T}\sum_{k=1}^{n}\mathbb{E}\big{[}\mathbb{E}[\boldsymbol{1}_{\{\alpha^{t}_{m_{k}}\neq w_{k}\}}|\hat{X}^{t}_{m_{k}}]\big{|}\Omega\big{]}=\sum_{t=1}^{T}\sum_{k=1}^{n}\mathbb{E}\big{[}\mathbb{P}\{\alpha^{t}_{m_{k}}\neq w_{k}|\hat{X}^{t}_{m_{k}}\}\big{|}\Omega\big{]}$ (46)
$=\sum_{t=1}^{T}\sum_{k=1}^{n}\mathbb{E}\big{[}1-\hat{X}^{t}_{m_{k}w_{k}}|\Omega\big{]}\leq nt_{n}+\sum_{t=t_{n}+1}^{T}\sum_{k=1}^{n}\mathbb{E}\big{[}1-\hat{X}^{t}_{m_{k}w_{k}}|\Omega\big{]}$ (47)
$\leq\tilde{O}\big{(}\frac{n^{3}}{c^{2}}\log\frac{1}{\delta}\big{)}+\sum_{t=t_{n}+1}^{T}\sum_{k=1}^{n}\big{(}1-\frac{\gamma^{t}}{n}-\frac{1-\gamma^{t}}{1+(n-1)e^{-\eta^{t}a_{k}t}}\big{)}$ (48)
$\leq\tilde{O}\big{(}\frac{n^{3}}{c^{2}}\log\frac{1}{\delta}\big{)}+\sum_{t=t_{n}+1}^{T}\sum_{k=1}^{n}\frac{(n-1)e^{-\eta^{t}a_{k}t}+\gamma_{t}}{1+(n-1)e^{-\eta^{t}a_{k}t}}$ (49)
$\leq\tilde{O}\big{(}\frac{n^{3}}{c^{2}}\log\frac{1}{\delta}\big{)}+n^{2}\sum_{t=1}^{T}e^{-\eta^{t}a_{n}t}+n\sum_{t=1}^{T}\gamma_{t}$ (50)
$\leq\tilde{O}\big{(}\frac{n^{3}}{c^{2}}\log\frac{1}{\delta}\big{)}+\frac{2n^{2}}{a_{n}^{2}}+nM\log^{2}T$ (51)
$=O\Big{(}\frac{n^{3}}{c^{2}}\log\frac{1}{\delta}+\frac{n^{2}}{c}\log\frac{1}{\delta}\log^{2}T\Big{)},$ (52)
where the third inequality uses Lemma 3, the fifth inequality uses the fact that $a_{n}:=\min_{k\in[n]}\{a_{k}\}$, and the last inequality is obtained by choosing $\gamma^{t}=M\frac{\log t}{t},\eta^{t}=\frac{1}{\sqrt{t}},M=\frac{4n}{c}\log\frac{1}{\delta}$, and noting that $c\leq a_{n}$.
Finally, since $\delta\in(0,1)$ can be chosen arbitrary, by choosing $\delta=\frac{1}{T}$, we can write
$\mathbb{E}[R(T)]$ =$=\mathbb{P}\{\Omega\}\mathbb{E}[R(T)|\Omega]+\mathbb{P}\{\bar{\Omega}\}\mathbb{E}[R(T)|\bar{\Omega}]$
$\leq(1-\delta)O\Big{(}\frac{n^{3}}{c^{2}}\log\frac{1}{\delta}+\frac{n^{2}}{c}\log\frac{1}{\delta}\log^{2}T\Big{)}+\delta n^{2}T$ (53)
$=O\Big{(}\frac{n^{2}}{c}\log^{3}T\Big{)}.$ (54)
@@
Remark 3
It is worth noting that the regret of Algorithm 1 is logarithmic in $T$, with a polynomial factor $\frac{n^{2}}{c}$ in terms of the problem parameters, thereby addressing a question raised in [2]. Furthermore, Algorithm 1 is both decentralized and uncoordinated, overcoming the limitations of the algorithm proposed in [2], which is notably more complex and requires phase-dependent coordination among men according to the Deferred-Acceptance algorithm by Gale and Shapley.
IV Exponential Local Convergence for General Matching Markets
@@
Please see Appendix A-C for the proof.  Q.E.D.
Next, we consider the following lemma that will be used to capture the local behavior of the dynamics of Algorithm 1 around a pure NE.
@@
$|v_{mw}(\hat{X}^{\tau})-v_{mw}(X^{\tau})|$ $=|\nabla v_{mw}(\theta)\cdot(\hat{X}^{\tau}-X^{\tau})|$
$\leq\|\nabla v_{mw}(\theta)\|_{\infty}\|\hat{X}^{\tau}-X^{\tau}\|_{1}$ (59)
$\leq n\|\hat{X}^{\tau}-X^{\tau}\|_{1}$ (60)
$=n\gamma^{\tau}\sum_{m}\|X_{m}^{\tau}-\frac{1}{n}\boldsymbol{1}\|_{1}\leq n^{2}\gamma^{\tau}.$ (61)
@@
$|\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})|\leq\frac{n}{\gamma^{\tau}},$
$\mathbb{E}[\big{(}\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})\big{)}^{2}|\mathcal{F}^{\tau-1}]\leq\mathbb{E}[(\hat{v}_{mw}^{\tau})^{2}|\mathcal{F}^{\tau-1}]\leq\frac{n}{\gamma^{\tau}}.$ (64)
@@
Let us define $A^{t-1}=\sum_{\tau=1}^{t-1}\big{(}v_{mw^{*}}(X^{\tau})-v_{mw}(X^{\tau})\big{)}$. Using (62) and (65), with probability at least $1-\delta$, we have
@@
In fact, there exist feasible choices for $\eta^{t}$ and $\gamma^{t}$ that satisfy the stepsize condition (55) in Theorem 4. For instance, if for some $\alpha,\beta\in(0,1),\ \alpha>\max\{1-\beta,\frac{1+\beta}{2}\}$, we choose
$\eta^{t}=\Theta\Big{(}\frac{1}{\ln(t/\sqrt{\delta})t^{\alpha}}\Big{)},\ \ \ \ \ \gamma^{t}=\Theta(\frac{1}{t^{\beta}}),$
then condition (55) is satisfied for any $t_{0}\geq(\frac{\alpha b}{c})^{1/(1-\alpha)}$, where $b=\ln(\frac{n^{2}}{2c})+3$. More specifically, by choosing $\gamma^{t}=\frac{1}{t^{1/3}}$ and $\eta^{t}=\frac{1}{t^{3/4}}$, the exponential convergence rate of Theorem 4 would be in the order of $\exp(-ct^{1/4})$ for any $t_{0}\geq(\frac{b}{c})^{4}$. We note that condition (55) provides only a sufficient condition to achieve local exponential convergence and we did not attempt to fully characterize all feasible choices of stepsize/mixing parameters. However, there could be other choices of stepsize that do not satisfy condition (55) while still achieving local exponential convergence.
@@
Under the same stepsize assumptions of Theorem 4, if Algorithm 1 converges to a limit point $X^{*}$ with positive probability, then $X^{*}$ must be a (possibly mixed) NE of the matching game.
@@
Please see Appendix A-E for the proof.  Q.E.D.
V Global Learning of Stable Matchings in General Markets
In the previous section, we showed that the decentralized and uncoordinated EXP algorithm converges locally to a stable matching at an exponential rate in general matching markets. This prompts the question of whether there is an uncoordinated and decentralized algorithm capable of globally learning a stable matching in general markets, regardless of the convergence rate. In this section, we answer this question affirmatively by introducing an alternative algorithm that leverages the weak acyclic property of the stable matching game. To that end, we first present several useful properties of the stable matching game, which will be instrumental in establishing our main convergence results.

### v2

@@ IV-B Performance of Algorithm 1 for Hierarchical Markets
$Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w}$ $=\sum_{\tau=t_{k-1}}^{t-1}\big{(}v_{m_{k}w_{k}}(\hat{X}^{\tau})-v_{m_{k}w}(\hat{X}^{\tau})\big{)}+\sum_{\tau=1}^{t_{k-1}-1}\big{(}v_{m_{k}w_{k}}(\hat{X}^{\tau})-v_{m_{k}w}(\hat{X}^{\tau})\big{)}$
$\geq\sum_{\tau=t_{k-1}}^{t-1}\Big{(}\mu_{m_{k}w_{k}}\prod_{\ell=1}^{k-1}(1-\hat{X}^{\tau}_{m_{\ell}w_{k}})-\mu_{m_{k}w}\Big{)}-t_{k-1}$ (41)
$\geq\sum_{\tau=t_{k-1}}^{t-1}\Big{(}\mu_{m_{k}w_{k}}\big{(}1-\sum_{\ell=1}^{k-1}\hat{X}^{\tau}_{m_{\ell}w_{k}}\big{)}-\mu_{m_{k}w}\Big{)}-t_{k-1}$ (42)
$\geq\sum_{\tau=t_{k-1}}^{t-1}\Big{(}(\mu_{m_{k}w_{k}}-\mu_{m_{k}w})-\mu_{m_{k}w_{k}}\sum_{\ell=1}^{k-1}\big{(}\frac{\gamma^{\tau}}{n}+\frac{(1-\gamma^{t})e^{-\eta^{\tau}a_{\ell}\tau}}{1+(n-1)e^{-\eta^{\tau}a_{\ell}\tau}}\big{)}\Big{)}-t_{k-1}$ (43)
$\geq\sum_{\tau=t_{k-1}}^{t-1}\Big{(}\Delta-\gamma^{\tau}-\sum_{\ell=1}^{k-1}e^{-\eta^{\tau}a_{\ell}\tau}\Big{)}-t_{k-1}$ (44)
$\geq\sum_{\tau=t_{k-1}}^{t-1}\Big{(}\Delta-\gamma^{\tau}-ne^{-\eta^{\tau}a_{k-1}\tau}\Big{)}-t_{k-1}.$ (45)
Thus, for any integer $t\geq t_{k-1}$ and the choice of parameters $\eta^{t}=\frac{1}{\sqrt{t}}$ and $\gamma^{t}=M\frac{\log t}{t}$, we have
$Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w}$ $\geq\sum_{\tau=t_{k-1}}^{t-1}\Big{(}\Delta-\gamma^{\tau}-ne^{-\eta^{\tau}a_{k-1}\tau}\Big{)}-t_{k-1}$
$\geq\Delta(t-t_{k-1})-M\log^{2}t-\frac{4n}{a^{2}_{k-1}}-t_{k-1}$ (46)
$\geq\Delta t-M\log^{2}t-\frac{4n}{a^{2}_{k-1}}-2t_{k-1}.$ (47)
Therefore, for any $t\geq\frac{4}{\Delta}(M\log^{2}(\frac{2M}{\Delta})+\frac{2n}{a^{2}_{k-1}}+t_{k-1})$, we have
$Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w}\geq\frac{\Delta}{2}t.$ (48)
@@
$t_{k}:=\frac{4}{\min\{\Delta,\mu_{m_{k}w_{k}}\}}\Big{(}2M\log^{2}(\frac{2M}{\min\{\Delta,\mu_{m_{k}w_{k}}\}})+\frac{2n}{a^{2}_{k-1}}+t_{k-1}\Big{)},$ (49)
@@
$\frac{X^{t}_{m_{k}w_{k}}}{X^{t}_{m_{k}w}}$ $=e^{\eta^{t}(\hat{L}^{t-1}_{m_{k}w_{k}}-\hat{L}^{t-1}_{m_{k}w})}\geq e^{\eta^{t}(Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w})-2c\sqrt{t}}$
$\geq e^{\eta^{t}2a_{k}t-2c\sqrt{t}}\geq e^{\eta^{t}2a_{k}t-\eta^{t}a_{k}t}=e^{\eta^{t}a_{k}t},$ (52)
@@
Finally, using relation (49) recursively, and noting that for any $k$ we have $\frac{4}{\min\{\Delta,\mu_{m_{k}w_{k}}\}}\leq\frac{1}{c}$ and $\frac{2}{a_{k}}\leq\frac{1}{c}$, one can obtain an explicit upper bound for $t_{n}$ as
$t_{n}$ $\leq\frac{1}{c}\Big{(}2M\log^{2}(\frac{M}{c})+\frac{n}{c^{2}}+t_{n-1}\Big{)}\leq\cdots$
$\leq\Big{(}\sum_{k=1}^{n-1}\frac{1}{c^{k}}\Big{)}\cdot\Big{(}2M\log^{2}(\frac{M}{c})+\frac{n}{c^{2}}+t_{1}\Big{)}$ (53)
$=O\Big{(}\frac{nM}{c^{n+1}}\log^{2}(\frac{M}{c})\Big{)}.$ (54)
Q.E.D.
@@
Assume that each man follows Algorithm 1 with $\eta^{t}=\frac{1}{\sqrt{t}}$ and $\gamma^{t}=M\frac{\log t}{t}$, where $M=\frac{4n}{c}\log T$ and $c=\frac{1}{8}\min_{k\in[n]}\{\Delta,\mu_{m_{k}w_{k}}\}$. Then, the expected regret of Algorithm 1 in hierarchical matching markets is at most $R(T)=\tilde{O}\Big{(}\frac{n^{3}}{c^{n+2}}\log T+\frac{n^{2}}{c}\log^{3}T\Big{)}$, where $\tilde{O}$ hides the $\log(\log T)$ terms.
@@
Using the definition of the expected regret, we can write
$\mathbb{E}[R(T)]$ $=\sum_{t=1}^{T}\mathbb{E}[\boldsymbol{1}_{\{\exists k:\alpha^{t}_{m_{k}}\neq w_{k}\}}]$ (55)
$\leq\sum_{t=1}^{T}\sum_{k=1}^{n}\mathbb{E}[\boldsymbol{1}_{\{\alpha^{t}_{m_{k}}\neq w_{k}\}}]$ (56)
$=\sum_{t=1}^{T}\sum_{k=1}^{n}\mathbb{E}\big{[}\mathbb{E}[\boldsymbol{1}_{\{\alpha^{t}_{m_{k}}\neq w_{k}\}}|\hat{X}^{t}_{m_{k}}]\big{]}$ (57)
$=\sum_{t=1}^{T}\sum_{k=1}^{n}\mathbb{E}\big{[}\mathbb{P}\{\alpha^{t}_{m_{k}}\neq w_{k}|\hat{X}^{t}_{m_{k}}\}\big{]}$ (58)
$=\sum_{t=1}^{T}\sum_{k=1}^{n}\mathbb{E}\big{[}1-\hat{X}^{t}_{m_{k}w_{k}}\big{]}$ (59)
$\leq nt_{n}+\sum_{t=t_{n}+1}^{T}\sum_{k=1}^{n}\mathbb{E}\big{[}1-\hat{X}^{t}_{m_{k}w_{k}}\big{]}$ (60)
On the other hand, by conditioning on the event $\Omega$, and using the fact that $\mathbb{P}\{\bar{\Omega}\}\leq\delta$, we have
$\mathbb{E}\big{[}1-\hat{X}^{t}_{m_{k}w_{k}}\big{]}$ =$=\mathbb{P}\{\Omega\}\mathbb{E}\big{[}1-\hat{X}^{t}_{m_{k}w_{k}}|\Omega\big{]}+\mathbb{P}\{\bar{\Omega}\}\mathbb{E}\big{[}1-\hat{X}^{t}_{m_{k}w_{k}}|\bar{\Omega}\big{]}$ (61)
$\leq\mathbb{E}\big{[}1-\hat{X}^{t}_{m_{k}w_{k}}|\Omega\big{]}+\delta$ (62)
Substituting (61) into (55) and using Lemma 3, we obtain
$\mathbb{E}[R(T)]$ $\leq nt_{n}+\sum_{t=t_{n}+1}^{T}\sum_{k=1}^{n}\mathbb{E}\big{[}1-\hat{X}^{t}_{m_{k}w_{k}}|\Omega\big{]}+\delta nT$ (63)
$\leq O\Big{(}\frac{n^{2}M}{c^{n+1}}\log^{2}(\frac{M}{c})\Big{)}+\sum_{t=t_{n}+1}^{T}\sum_{k=1}^{n}\big{(}1-\frac{\gamma^{t}}{n}-\frac{1-\gamma^{t}}{1+(n-1)e^{-\eta^{t}a_{k}t}}\big{)}+\delta nT$ (64)
$\leq O\Big{(}\frac{n^{2}M}{c^{n+1}}\log^{2}(\frac{M}{c})\Big{)}+\sum_{t=t_{n}+1}^{T}\sum_{k=1}^{n}\frac{(n-1)e^{-\eta^{t}a_{k}t}+\gamma_{t}}{1+(n-1)e^{-\eta^{t}a_{k}t}}+\delta nT$ (65)
$\leq O\Big{(}\frac{n^{2}M}{c^{n+1}}\log^{2}(\frac{M}{c})\Big{)}+n^{2}\sum_{t=1}^{T}e^{-\eta^{t}a_{n}t}+n\sum_{t=1}^{T}\gamma_{t}+\delta nT$ (66)
$\leq O\Big{(}\frac{n^{2}M}{c^{n+1}}\log^{2}(\frac{M}{c})\Big{)}+\frac{2n^{2}}{c^{2}}+nM\log^{2}T+\delta nT$ (67)
$=O\Big{(}\frac{n^{2}M}{c^{n+1}}\log^{2}(\frac{M}{c})+nM\log^{2}T+\delta nT\Big{)},$ (68)
where the fourth inequality uses the fact that $a_{n}:=\min_{k\in[n]}\{a_{k}\}$, and the last inequality is obtained by choosing $\gamma^{t}=M\frac{\log t}{t},\eta^{t}=\frac{1}{\sqrt{t}}$, and noting that $c\leq a_{n}$. Finally, since $\delta\in(0,1)$ can be chosen arbitrarily, by choosing $\delta=\frac{1}{T}$ and noting that $M=\frac{4n}{c}\log\frac{1}{\delta}=\frac{4n}{c}\log T$, we can write
$\mathbb{E}[R(T)]$ $\leq O\Big{(}\frac{n^{2}M}{c^{n+1}}\log^{2}(\frac{M}{c})+nM\log^{2}T+\delta nT\Big{)}$
$=O\Big{(}\frac{n^{3}}{c^{n+2}}\log T\cdot\log^{2}(\log T)+\frac{n^{2}}{c}\log^{3}T\Big{)}.$ (69)
@@
V Exponential Local Convergence for General Matching Markets
@@
Please see Appendix B-B for the proof. Q.E.D.
Next, we consider the following lemma that will be used to capture the local behavior of the dynamics of Algorithm 1 around a pure NE. The proof of this lemma can be found in Appendix B-C.
@@
$|v_{mw}(\hat{X}^{\tau})-v_{mw}(X^{\tau})|$ $=|\nabla v_{mw}(\theta)\cdot(\hat{X}^{\tau}-X^{\tau})|\leq\|\nabla v_{mw}(\theta)\|_{\infty}\|\hat{X}^{\tau}-X^{\tau}\|_{1}$
$\leq n\|\hat{X}^{\tau}-X^{\tau}\|_{1}=n\gamma^{\tau}\sum_{m}\|X_{m}^{\tau}-\frac{1}{n}\boldsymbol{1}\|_{1}\leq n^{2}\gamma^{\tau}.$ (74)
@@
$|\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})|\leq\frac{n}{\gamma^{\tau}},\ \ \ \ \mathbb{E}[\big{(}\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})\big{)}^{2}|\mathcal{F}^{\tau-1}]\leq\mathbb{E}[(\hat{v}_{mw}^{\tau})^{2}|\mathcal{F}^{\tau-1}]\leq\frac{n}{\gamma^{\tau}}.$
@@
Let $A^{t-1}=\sum_{\tau=1}^{t-1}\big{(}v_{mw^{*}}(X^{\tau})-v_{mw}(X^{\tau})\big{)}$. Using (75) and (77), with probability at least $1-\delta$, we have
@@
In fact, there exist feasible choices for $\eta^{t}$ and $\gamma^{t}$ that satisfy the stepsize condition (70) in Theorem 4. For instance, if for some $\alpha,\beta\in(0,1),\ \alpha>\max\{1-\beta,\frac{1+\beta}{2}\}$, we choose $\eta^{t}=\Theta\big{(}\frac{1}{\ln(t/\sqrt{\delta})t^{\alpha}}\big{)}$ and $\gamma^{t}=\Theta(\frac{1}{t^{\beta}})$, then condition (70) is satisfied for any $t_{0}\geq(\frac{\alpha b}{c})^{1/(1-\alpha)}$, where $b=\ln(\frac{n^{2}}{2c})+3$. More specifically, by choosing $\gamma^{t}=\frac{1}{t^{1/3}}$ and $\eta^{t}=\frac{1}{t^{3/4}}$, the exponential convergence rate of Theorem 4 would be in the order of $\exp(-ct^{1/4})$ for any $t_{0}\geq(\frac{b}{c})^{4}$. We note that condition (70) provides only a sufficient condition to achieve local exponential convergence and we did not attempt to fully characterize all feasible choices of stepsize/mixing parameters. However, there could be other choices of stepsize that do not satisfy condition (70) while still achieving local exponential convergence.
@@
Under the same stepsize assumptions of Theorem 4, if Algorithm 1 converges to a limit point $X^{*}$ with positive probability, then $X^{*}$ must be a NE of the stable matching game.
@@
For any $w^{\prime}\neq w$, by taking limit from the above relation as $t\to\infty$, we get $\limsup_{t\to\infty}\eta^{t}A^{t-1}\leq\ln(\frac{X^{*}_{mw^{\prime}}}{X^{*}_{mw}})+3\leq O(1)$, where if $X^{*}_{mw^{\prime}}=0$, by convention we define $\ln(\frac{X^{*}_{mw^{\prime}}}{X^{*}_{mw}})=-\infty$. Now let $\kappa:=v_{mw^{\prime}}(X^{*})-v_{mw}(X^{*})$. We claim $\kappa\leq 0$. Otherwise, if $\kappa>0$, then for any sufficiently large $t>\bar{t}$ we must have $A^{t-1}>t\kappa/2$. Given the choice of stepsize, we obtain $\lim_{t\to\infty}\eta^{t}A^{t-1}=\lim_{t\to\infty}t\eta^{t}\kappa/2=\infty$, which contradicts the fact that $\limsup_{t\to\infty}\eta^{t}A^{t-1}\leq O(1)$. Thus, $\kappa\leq 0$, which shows that if $X^{*}_{mw}>0$, then $v_{mw}(X^{*})=\max_{w^{\prime}}v_{mw^{\prime}}(X^{*})$, i.e., $X^{*}$ is a NE. Q.E.D.
VI Global Learning of Stable Matchings in General Markets
In the previous section, we showed that the decentralized and uncoordinated EXP algorithm converges locally to a stable matching at an exponential rate in general matching markets. This prompts the question of whether there is an uncoordinated and decentralized algorithm capable of globally learning a stable matching in general markets, regardless of the convergence rate. In this section, we raiser this question affirmatively by introducing an alternative algorithm that leverages the weak acyclic property of the stable matching game. To that end, we first present several useful properties of the stable matching game, which will be instrumental in establishing our main convergence results.